2. Craps is a game played by rolling two fair dice. To play one round of this game, the player rolls the dice and the outcome is determined by the following rules: If the total number of dots is 7 or 11 (a natural), then the player wins. If the total number of dots is 2, 3, or 12 Ccraps), then the player loses. If the total number of dots is 4, 5, 6,8,9, or 10, then this number is called the players point value and the player must continue rolling the dice until they roll the point value again or they roll a total of 7. If they roll the point value before a total of 7 appears, the player wins. If they roll a total of 7 before the point value appears, the player loses. What is the probability that a person playing one round of this game will win?
Answer #1

Probability that a person playing one round of the game wins =

P(the person wins with first outcome 7) + P(the person wins with first outcome 11) + P(the person wins with first outcome 4) + P(the person wins with first outcome 5) + P(the person wins with first outcome 6) + P(the person wins with first outcome 8) + P(the person wins with first outcome 9) + P(the person wins with first outcome 10)

.

P(first outcome is 7) = P(1,6)+P(6,1)+P(2,5)+P(5,2)+P(3,4)+P(4,3) = 6/36 = 1/6

P(first outcome is 11) = P(5,6)+P(6,5) = 2/36 = 1/18

P(first outcome is 4) = P(1,3)+P(3,1)+P(2,2) = 3/36 = 1/12

P(first outcome is 5) = P(1,4)+P(4,1)+P(2,3)+P(3,2) = 4/36 = 1/9

P(first outcome is 6) = P(1,5)+P(5,1)+P(2,4)+P(4,2)+P(3,3) = 5/36

P(first outcome is 8) = P(2,6)+P(6,2)+P(3,5)+P(5,3)+P(4,4) = 5/36

P(first outcome is 9) = P(3,6)+P(6,3)+P(4,5)+P(5,4) = 4/36 = 1/9

P(first outcome is 10) = P(4,6)+P(6,4)+P(5,5) = 3/36 = 1/12

.

P(the person wins with first outcome 7) = P(first outcome is 7) = 1/6

P(the person wins with first outcome 11) = P(first outcome is 11) = 1/18

P(the person wins with first outcome 4)

= P(first outcome is 4 , second outcome is 4) + P(first outcome is 4 , second outcome is not 7 , third outcome is 4) + P(first outcome is 4 , second and third outcome is not 4 , fourth outcome is 4) + ........

= (1/12)(1/12) + (1/12)(1-1/6)(1/12) + (1/12)(1-1/6)2(1/12) + .........

= (1/12)2 + (1/12)2(5/6) + (1/12)2(5/6)2 + .........

=11/12/6) = 24

P(the person wins with first outcome 5)

= P(first outcome is 5 , second outcome is 5) + P(first outcome is 5 , second outcome is not 7 , third outcome is 5) + P(first outcome is , second and third outcome is not 4 , fourth outcome is 5) + ........

= (1/9)(1/9) + (1/9)(1-1/6)(1/9) + (1/9)(1-1/6)2(1/9) + .........

= (1/9)2 + (1/9)2(5/6) + (1/9)2(5/6)2 + .........

(1/9/6) = 27

P(the person wins with first outcome 6)

= P(first outcome is 6 , second outcome is 6) + P(first outcome is 6 , second outcome is not 7 , third outcome is 6) + P(first outcome is 6 , second and third outcome is not 4 , fourth outcome is 6) + ........

= (5/36)(5/36) + (5/36)(1-1/6)(5/36) + (5/36)(1-1/6)2(5/36) + .........

= (5/36)2 + (5/36)2(5/6) + (5/36)2(5/6)2 + .........

(5/36) 25 1 - (5/6) 216

P(the person wins with first outcome 8)

= P(first outcome is 8 , second outcome is 8) + P(first outcome is 8 , second outcome is not 7 , third outcome is 8) + P(first outcome is 8 , second and third outcome is not 4 , fourth outcome is 8) + ........

= (5/36)(5/36) + (5/36)(1-1/6)(5/36) + (5/36)(1-1/6)2(5/36) + .........

= (5/36)2 + (5/36)2(5/6) + (5/36)2(5/6)2 + .........

(5/36) 25 1 - (5/6) 216

P(the person wins with first outcome 9)

= P(first outcome is 9 , second outcome is 9) + P(first outcome is 9 , second outcome is not 7 , third outcome is 9) + P(first outcome is 9 , second and third outcome is not 4 , fourth outcome is 9) + ........

= (1/9)(1/9) + (1/9)(1-1/6)(1/9) + (1/9)(1-1/6)2(1/9) + .........

= (1/9)2 + (1/9)2(5/6) + (1/9)2(5/6)2 + .........

(1/9/6) = 27

P(the person wins with first outcome 10)

= P(first outcome is 10 , second outcome is 10) + P(first outcome is 10 , second outcome is not 7 , third outcome is 10) + P(first outcome is 10 , second and third outcome is not 4 , fourth outcome is 10) + ........

= (1/12)(1/12) + (1/12)(1-1/6)(1/12) + (1/12)(1-1/6)2(1/12) + .........

= (1/12)2 + (1/12)2(5/6) + (1/12)2(5/6)2 + .........

=11/12/6) = 24

.

herefore  Probability that a person playing one round of the game wins

1 12252521 37 6 18 24 27 216 216 27 24 54